OpenStudy (anonymous):

MEDAL FOR ANSWER!!!!!!!!! Which of the following is an equivalent form of the compound inequality -22 > -5x - 7 -3? A)-5x - 7 < -22 and -5x - 7 -3 B)-5x - 7 < -22 and -5x - 7 -3 C)-5x - 7 > -22 and -5x - 7 -3 D)-5x > -22 and -7 -3

OpenStudy (mertsj):

Something missing in your post.

OpenStudy (anonymous):

your right

OpenStudy (mertsj):

I know.

OpenStudy (anonymous):

-22 > -5x - 7 greater than or equal to -3?

OpenStudy (mertsj):

\[-22>-5x-7\ge-3\] \[-22+7>-5x \ge4\] \[-15>-5x \ge4\] \[\frac{-15}{-5}<\frac{-5x}{-5}\le\frac{4}{-5}\] \[3<x \le\frac{-4}{5}\]

OpenStudy (mertsj):

Your answer choices are missing inequality symbols.

OpenStudy (mertsj):

Equivalent statements are: \[-5x-7<-22 and -5x-7\ge-3\]

OpenStudy (anonymous):

because you cant copy/past a greater then or equal too symbol.

OpenStudy (mertsj):

That is a problem.

OpenStudy (anonymous):

yea and thank you and can you teach me how to graph really quick

Directrix (directrix):

-22 > -5x - 7 greater than or equal to -3 when pulled apart means -22 > -5x -7 AND -5x - 7 ≥ -3 -5x -7 < -22 AND -5x - 7 ≥ -3 You forgot to put the inequality symbols in the last part of each answer option. Please add those.

OpenStudy (mertsj):

What do you want to graph?

OpenStudy (anonymous):

Which of the following is the correct graph of the compound inequality n > 5 and n < 3?

OpenStudy (mertsj):

First we see that we will have open circles because it is > and < and greater than or equal to or less than or equal to.

OpenStudy (mertsj):

So put an open circle on 5 and a line with an arrow going to the right since it says that n is greater than 5.

OpenStudy (mertsj):

Then put an open circle on 3 and a line with an arrow going to the left since it says that n is less than 3.

OpenStudy (mertsj):



OpenStudy (anonymous):

Which of the following is the correct graph of the compound inequality 4p + 1 > -7 or 6p + 3 < 33?

OpenStudy (mertsj):

You first must solve both of those for p. Can you do that?

OpenStudy (anonymous):

p+1=-3??

OpenStudy (mertsj):